OpenStudy (anonymous):

Will fan+medal. prove the following identity: (1/tan A)+ (cos(90-A))/(cos A+1)=cosec A

OpenStudy (anonymous):

\[\frac{ 1 }{ \tan A}+\frac{ \cos(90-A) }{ \cos A+1 }= cosecA\]